OpenStudy (anonymous):

Evaluate using a substitution: lim x->2+ of arctan((1/(x-2))

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2+} \arctan (\frac{ 1 }{ x-2 })\]

OpenStudy (lgbasallote):

so what happens if you substitute x with 2?

OpenStudy (anonymous):

The number inside the parenthesis becomes undefined (1/0)

OpenStudy (lgbasallote):

and what is the arctangent of undefined?

OpenStudy (lgbasallote):

okay... here's a hint: \[\tan = \frac{\sin}{\cos}\] that will become undefined if cosine is 0... so what value of x will make cos x = 0?

OpenStudy (anonymous):

pi/2 makes cos x =0

OpenStudy (lgbasallote):

right. \[\tan (\frac \pi 2) = \frac{\sin (\frac \pi 2)}{\cos ( \frac \pi 2)} \implies \tan (\frac \pi 2) = \frac 10 \implies \text{undefined}\] do you get it now?

OpenStudy (anonymous):

I understand that but how does this help if its arctan?

OpenStudy (lgbasallote):

if the tangent of pi/2 is undefined...then what do you think is the arctangent of undefined?

OpenStudy (anonymous):

OHHH its pi/2

OpenStudy (lgbasallote):

indeed

OpenStudy (anonymous):

But like is that how i wud show work? Like on an exam instead of reasoning with teh tangent is tehre a way to show taht its indeed pi/2?

OpenStudy (lgbasallote):

actually..you don't ahve to show the solution of the tangent...arctangent of (1/0) is already automatic

OpenStudy (lgbasallote):

i just showed you the solution of the tangent for you to see

OpenStudy (anonymous):

B/C im supposed to evaluate using subsition. So what shud I write on paper as my work then?

OpenStudy (anonymous):

Wait is like arctan of undefined always pi/2 b/c thats just smthg we need to memorize just like sin(90) = 1

OpenStudy (anonymous):

THANKS NEVER MIND I GET IT :D

OpenStudy (lgbasallote):

arctan (undefined) is not automatic 90 degrees by the way... it's 90 degrees in this case ebcause the numerator is 1 and 1 can only be got with sin (90)